Solve using the addition principle x+1/13=5/13. The solution is X=??
Answer by ewatrrr(24785)


Hi
x+1/13=5/13 |Adding -1/13 to both sides of the equation
x+1/13 -1/13 =5/13 -1/13
x = 4/13